AB INITIO CALCULATION OF THE ELECTRONIC TRANSITIONS OF Er^{3+} DOPED INTO GaN

\underline{YANG YANG} and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

As a potential application in optical communications, the 1.54 μm photoluminescence (^{4}I_{13/2} →^{4}I_{15/2} intra-4f shell emissions) of Er^{3+} doped in the semiconductor GaN has been studied extensively since 1983. Few ab initio calculations touched this system before because of the difficulties dealing with the large number of electrons and the significant relativistic effects related to Er^{3+}. Based on relativistic effective core potentials, spin-orbit configuration interaction calculations have been performed on this system. The study of the wave function character and energy of both the ground and excited states of ErN_{4}^{9−} plus the computation of the transition moments between the states gives a better understanding of the experimental results.